# Enhancing Test Time Adaptation with Few-shot Guidance

### v1

@@ Abstract
Deep neural networks often encounter significant performance drops while facing with domain shifts between training (source) and test (target) data. To address this issue, Test Time Adaptation (TTA) methods have been proposed to adapt pre-trained source model to handle out-of-distribution streaming target data. Although these methods offer some relief, they lack a reliable mechanism for domain shift correction, which can often be erratic in real-world applications. In response, we develop Few-Shot Test Time Adaptation (FS-TTA), a novel and practical setting that utilizes a few-shot support set on top of TTA. Adhering to the principle of few inputs, big gains, FS-TTA reduces blind exploration in unseen target domains. Furthermore, we propose a two-stage framework to tackle FS-TTA, including (i) fine-tuning the pre-trained source model with few-shot support set, along with using feature diversity augmentation module to avoid overfitting, (ii) implementing test time adaptation based on prototype memory bank guidance to produce high quality pseudo-label for model adaptation. Through extensive experiments on three cross-domain classification benchmarks, we demonstrate the superior performance and reliability of our FS-TTA and framework.
Introduction
In recent years, deep neural networks have exhibited remarkable capabilities in representation learning. However, their performance relies heavily on the assumption that the distributions of training (source) and test (target) data are identical (Long et al. 2015; Ganin and Lempitsky 2015; Li et al. 2017). In real-world deployment, such a distribution shift is inevitable, as it is practically impossible to collect and annotate data for all possible environments in advance of training. Besides, this distribution shift can significantly degrade the performance of the deployed source model.
To address the aforementioned issues, numerous studies have proposed solutions via domain adaptation (Long et al. 2015; Tzeng et al. 2017; Zhang et al. 2019; Xiao and Zhang 2021; Xin et al. 2023) and domain generalization (Volpi et al. 2018; Zhou et al. 2021; Kim et al. 2021; Sicilia, Zhao, and Hwang 2023). While these approaches have demonstrated impressive performance gains on realistic benchmarks, a considerable gap remains between their problem settings and practical application scenarios. Domain adaptation relies on the impractical assumption that target domain data are available and participate in the source training process. In contrast, domain generalization aims to directly enhance the generalization of the source model without exploring the target domain data, even if they can be obtained during the test time.
@@ Introduction
Figure 1: Test Time Adaptation (TTA) vs. Few-Shot Test Time Adaptation (FS-TTA). FS-TTA utilizes labeled few-shot target data in addition to TTA. The results for TTA are derived from the performance of TENT (Wang et al. 2021) on the OfficeHome (Venkateswara et al. 2017).
@@
In order to overcome these limitations of domain adaptation/generalization and protect the privacy of the source data, TENT (Wang et al. 2021) introduces fully test time adaptation (TTA). TTA aims to adapt a pre-trained source model to the target domain using input mini-batch data during the test time, without relying on source data or supervision. TTA is particularly focused on an online setting, where the model must adapt and make predictions immediately upon receiving each batch of potentially non-independent and identically distributed (non-i.i.d.) target samples. To serve this purpose, TENT employs test-time entropy minimization to reduce the generalization error on shifted target data. Additionally, extensive research has sought to improve TTA through various approaches such as pseudo-labeling (Iwasawa and Matsuo 2021; Wang, Zhang et al. 2023), consistency regularization (Boudiaf et al. 2022), and anti-forgetting regularization (Niu et al. 2022). While these methods can perform model adaptation during the test time, they encounter three primary challenges:
@@
1) Domain shift correction: The certainty of TTA methods in addressing domain shifts effectively without utilizing target labels is questionable. The t-SNE visualization in Fig. 5 clearly illustrates this point, where we observe that the feature distribution exhibits negligible change following the adaptation process with TENT. This suggests that TTA methods may struggle to effectively adjust to new domain characteristics in the complete absence of target labels, which could provide essential guidance for adaptation.
@@
2) Generalizability: The effectiveness of TTA methods varies across different scenarios. In some cases, they might even underperform compared to the pre-trained source model without any adaptation, as illustrated in Fig. 1 (Source Model vs. TENT). This variability indicates that the generalization performance of TTA methods is not particularly strong and can be influenced by various factors, including the domain shift and the specific characteristics of the model and dataset involved.
@@
Figure 2: The performance impact of one-shot.
The fundamental reason for these challenges is the blind exploration of the target domain without any supervision information when domain shift exists. Actually, in practical scenarios, it’s quite feasible to obtain few-shot labeled samples from the target domain. Consequently, this leads us to ask: If given little target domain supervisory information, could the adaptation performance be improved?
To answer this question, we test the one-shot situation, as shown in Fig. 2. Specifically, we use one sample per class to fine-tune the source model with cross-entropy loss. We find that the performance is easily improved compared to TENT, which shows that little supervision information can be more effective than a large amount of unsupervised information.
Based on the aforementioned findings, we introduce Few-Shot Test Time Adaptation (FS-TTA), encapsulating the concept of few inputs, big gains. As illustrated in Fig. 1, by integrating a few-shot support set from the target domain prior to model adaptation, FS-TTA effectively reduces domain shift while retaining the source-free and online characteristics inherent to TTA. It is noted that such setting is particularly beneficial in scenarios where precision and reliability are paramount, such as medical image analysis and autonomous driving, even spending a few extra labeling costs.
@@
  • New research direction: We highlight the setting of Few-Shot Test Time Adaptation (FS-TTA), in which an additional few-shot support set is available prior to test time adaptation. Thanks to the few-shot samples, domain shift can be reduced more effectively.
  • Novel framework: We propose a carefully crafted framework to tackle FS-TTA, including fine-tuning the pre-trained source model with feature diversity augmentation module and performing test time adaptation via high-quality pseudo-labeled samples in self-training manner.
  • Superior performance: We provide empirical results on various cross-domain classification benchmarks to demonstrate the superiority of our framework. In comparison to the state-of-the-art method in TTA, our method achieves improvements of 2.0% on PACS, 7.8% on OfficeHome, and 3.9% on DomainNet.
Related Work
Test Time Adaptation
Test Time Adaptation (TTA) strives to adapt the pre-trained source model during the test time to alleviate distribution shifts. Previous TTA methods address distribution alignment by incorporating self-supervised tasks (Sun et al. 2020). Although effective, TTA necessitates access to training samples and modifies the training process. To surmount this challenge, TENT (Wang et al. 2021) introduces the concept of fully test-time adaptation, utilizing only target data. TENT adapts the batch normalization layer by minimizing the entropy of model predictions. Distinct from TENT, recent works (Schneider et al. 2020; Nado et al. 2020) estimate the mean and variance of the activation in each batch normalization on the forthcoming test mini-batch. Consequently, LAME (Boudiaf et al. 2022) and EATA (Niu et al. 2022) concentrate on tackling the issue of catastrophic forgetting as the source model undergoes continuous updates. TSD (Wang, Zhang et al. 2023) integrates self-training into TTA, selecting high-quality test samples to fit the model.
@@
[…]
Comparisons with Other Settings
@@
We compare Few-Shot Test Time Adaptation (FS-TTA) with similar problem settings (details are in the appendix), as illustrated in Tab. 1.
@@
  • Compared with Domain Generalization, FS-TTA eliminates the need for access to source data, thereby safeguarding the privacy of the source data. Moreover, FS-TTA facilitates adaptation to the downstream target domain via model parameter updates.
  • Compared with Source-Free Domain Adaptation, FS-TTA obviates the need to acquire all target domain data simultaneously and can dynamically update the model online, contingent upon incoming mini-batches.
@@
  • Compared with Few-Shot Transfer Learning, FS-TTA can not only utilize a small number of target samples, but can also further update the model using online mini-batch target data during the test time.
@@
  • Compared with Test Time Adaptation, FS-TTA capitalizes on an auxiliary small set of samples from the target, allowing the pre-trained source model to adapt more swiftly and proficiently to the target domain. Furthermore, FS-TTA demonstrates exceptional performance in managing circumstances involving substantial domain shifts.
Method
Problem Setting
Considering a typical scenario where a source model $f_{\theta_{s}}$, is equipped with parameters $\theta_{s}$ and trained on source datasets $\mathcal{D}_{s_{1}},\mathcal{D}_{s_{2}},...,\mathcal{D}_{s_{n}}$, our goal is to adapt this pre-trained source model to a target domain $D_{t}$ without accessing source data. We postulate the existence of a small, labeled support set $S=\{(s_{i},y_{i})\}$ derived from $D_{t}$, where $s_{i}$ represents the image, and $y_{i}$ is its corresponding label. During the test time, mini-batches unlabeled target samples online arrive. Few-Shot Test Time Adaptation (FS-TTA) strives to effectively adapt the pre-trained source model $f_{\theta_{s}}$, utilizing the labeled support set $S$ and unlabeled online mini-batch to tackle the domain shift challenge. Notably, the support set $S$ could be offline obtained before the test time.
Stage I: Model Adaptation via Fine-Tuning
To significantly and swiftly enhance the initialization performance of the pre-trained source model in the target domain and minimize domain shifts, we design to fine-tune the pre-trained source model using the few-shot support set. Given the limited number of samples per class, there is a potential risk of overfitting during the fine-tuning process. To mitigate this, we introduce the Feature Diversity Augmentation (FDA) module, which generates new features by mixing statistics. Ultimately, we use a supervised classification loss to fine-tune the pre-trained source model. This entire procedure is illustrated in Stage I of Fig. 3.
@@
Feature Diversity Augmentation.
Prior research (Zhou et al. 2021) has demonstrated a significant association between feature statistics and image style, which is intricately linked to data distribution within the field of computer vision. Inspired by their discoveries, we propose Feature Diversity Augmentation (FDA), which enhances feature diversity and mitigates the risk of overfitting during fine-tuning.
FDA is incorporated between layers (blocks) in the pre-trained source backbone, as depicted in Fig. 3. More specifically, FDA mixes the feature statistics of two random samples to generate new features. The computations within FDA module can be summarized in three steps. Firstly, given two feature maps $f_{i}$ and $f_{j}$ from the support set, we compute their feature statistics ($\mu_{i},\sigma_{i}$) and ($\mu_{j},\sigma_{j}$). Secondly, FDA generates the mixtures of feature statistics:
$\boldsymbol{\gamma}_{\text{mix }}=\lambda\sigma_{i}+(1-\lambda)\sigma_{j},$ (1)
$\boldsymbol{\beta}_{\text{mix }}=\lambda\mu_{i}+(1-\lambda)\mu_{j}.$ (2)

### v2

@@ Abstract
Deep neural networks often experience significant performance degradation when encountering domain shifts between training (source) and test (target) data. To mitigate this issue, Test Time Adaptation (TTA) methods have been introduced to adapt a pre-trained source model for handling out-of-distribution streaming target data. While these methods provide some improvement, they often lack a dependable mechanism for domain shift correction, which can be unpredictable in real-world applications. To address this challenge, we propose the Few-Shot Test Time Adaptation (FS-TTA), a novel and practical setting that builds upon TTA by leveraging a few-shot support set. Following the principle of few inputs, big gains, FS-TTA effectively reduces the uncertainty of blind adaptation in previously unseen target domains. Furthermore, we introduce a two-stage framework to effectively tackle FS-TTA. The first stage involves fine-tuning the pre-trained source model using the few-shot support set while incorporating a feature diversity augmentation module to prevent overfitting. The second stage utilizes test time adaptation guided by a prototype memory bank, which facilitates the generation of high-quality pseudo-labels for model adaptation. Extensive experiments conducted on three cross-domain classification benchmarks validate the effectiveness, reliability, and superior performance of our proposed FS-TTA and framework.
keywords:
Test Time Adaptation, Few-Shot, Domain Shift, Transfer Learning.
@@
Figure 1: Test Time Adaptation (TTA) vs. Few-Shot Test Time Adaptation (FS-TTA). FS-TTA incorporates a small number of labeled target samples, which can be easily collected offline before deployment with minimal annotation effort, in addition to the unlabeled target data used in TTA. The results for TTA are based on the performance of TENT (Wang et al, 2021) on the OfficeHome (Venkateswara et al, 2017).
1 Introduction
In recent years, deep neural networks have exhibited remarkable capabilities in representation learning. However, their performance relies heavily on the assumption that the distributions of training (source) and test (target) data are identical (Long et al, 2015; Ganin and Lempitsky, 2015; Li et al, 2017). In real-world deployment, such a distribution shift is inevitable, as it is practically impossible to collect and annotate data for all possible environments in advance of training. Besides, this distribution shift can significantly degrade the performance of the deployed source model.
To address the aforementioned issues, numerous studies have proposed solutions via domain adaptation (Long et al, 2015; Tzeng et al, 2017; Zhang et al, 2019; Xiao and Zhang, 2021; Xin et al, 2023) and domain generalization (Volpi et al, 2018; Zhou et al, 2021; Kim et al, 2021; Sicilia et al, 2023). While these approaches have demonstrated impressive performance gains on realistic benchmarks, a considerable gap remains between their problem settings and practical application scenarios. Domain adaptation relies on the impractical assumption that target domain data are available and participate in the source training process. In contrast, domain generalization aims to directly enhance the generalization of the source model without exploring the target domain data, even if they can be obtained during the test time.
@@ 1 Introduction
In order to overcome these limitations of domain adaptation/generalization and protect the privacy of the source data, TENT (Wang et al, 2021) introduces fully test time adaptation (TTA). TTA aims to adapt a pre-trained source model to the target domain using input mini-batch data during the test time, without relying on source data or supervision. TTA is particularly focused on an online setting, where the model must adapt and make predictions immediately upon receiving each batch of potentially non-independent and identically distributed (non-i.i.d.) target samples. To serve this purpose, TENT employs test-time entropy minimization to reduce the generalization error on shifted target data. Additionally, extensive research has sought to improve TTA through various approaches such as pseudo-labeling (Iwasawa and Matsuo, 2021; Wang et al, 2023), consistency regularization (Boudiaf et al, 2022), and anti-forgetting regularization (Niu et al, 2022). While these methods can perform model adaptation during the test time, they encounter three primary challenges:
@@
1) Domain shift correction: The certainty of TTA methods in addressing domain shifts effectively without utilizing target labels is questionable. The t-SNE visualization in Figure 7 clearly illustrates this point, where we observe that the feature distribution exhibits negligible change following the adaptation process with TENT. This suggests that TTA methods may struggle to effectively adjust to new domain characteristics in the complete absence of target labels, which could provide essential guidance for adaptation.
@@
2) Generalizability: The effectiveness of TTA methods varies across different scenarios. In some cases, they might even underperform compared to the pre-trained source model without any adaptation, as illustrated in Figure 1 (Source Model vs. TENT). This variability indicates that the generalization performance of TTA methods is not particularly strong and can be influenced by various factors, including the domain shift and the specific characteristics of the model and dataset involved.
@@
Figure 2: Performance comparison across different adaptation strategies on OfficeHome. One-shot fine-tuning with a single labeled sample per class already surpasses TENT, showing the effectiveness of minimal supervision.
The fundamental reason for these challenges lies in the blind exploration of the target domain without any supervisory signal under domain shift. In practical scenarios, however, it is often feasible to obtain a small number of labeled target samples with minimal annotation cost. For instance, a few representative examples can be labeled offline before deployment by domain experts or system users. This leads us to ask: If given limited supervisory information from the target domain, could the adaptation performance be improved?
To answer this question, we test the one-shot situation, as shown in Figure 2. Specifically, we use one sample per class to fine-tune the source model with cross-entropy loss. We find that the performance is easily improved compared to TENT, which shows that little supervision information can be more effective than a large amount of unsupervised information.
Based on the aforementioned findings, we introduce the Few-Shot Test Time Adaptation (FS-TTA), encapsulating the concept of few inputs, big gains. As illustrated in Figure 1, by integrating a few-shot support set from the target domain prior to model adaptation, FS-TTA effectively reduces domain shift while retaining the source-free and online characteristics inherent to TTA. It is noted that such setting is particularly beneficial in scenarios where precision and reliability are paramount, such as medical image analysis and autonomous driving, even spending a few extra labeling costs.
@@
  • Emerging research direction: We highlight the setting of the Few-Shot Test Time Adaptation (FS-TTA), where a limited few-shot support set is available prior to test-time adaptation. Leveraging these few-shot samples enables more effective mitigation of domain shift.
  • Innovative framework: We propose a meticulously designed framework to address FS-TTA, incorporating fine-tuning of the pre-trained source model with a Feature Diversity Augmentation (FDA) module and performing test-time adaptation through high-quality pseudo-labeled samples in a self-training manner.
  • State-of-the-Art Performance: Extensive empirical evaluations on multiple cross-domain classification benchmarks validate the effectiveness of our framework. Compared to the current state-of-the-art TTA methods, our approach achieves performance gains of 2.0% on PACS, 7.8% on OfficeHome, and 3.9% on DomainNet.
@@
Domain Generalization $L^{s_{1}},\dots,L^{s_{N}}$ 0
Source-Free Domain Adaptation Pre-trained model on $L^{s_{1}},\dots,L^{s_{N}}$ Entire $U^{t}$ $\|U^{t}\|$
Few-Shot Transfer Learning Pre-trained model on $L^{s_{1}},\dots,L^{s_{N}}$ Few-shot support set $L^{spt}\subset L^{t}$ $k\times\text{C}$
@@
2 Related Work
2.1 Domain Generalization
Domain Generalization (DG) aims to train models on multiple related but distinct source domains to ensure effective performance on unseen target domains. To enhance robustness, DG techniques often employ strategies such as data augmentation (Huang et al, 2021; Volpi et al, 2018) and data generation (Zhou et al, 2021; Robey et al, 2021) to introduce greater diversity during training. Other prevalent approaches leverage representation learning to extract domain-invariant features. This includes kernel-based methods (Li et al, 2018b) that project data into a shared feature space, domain adversarial learning (Sicilia et al, 2023) that aligns distributions via adversarial objectives, and invariant risk minimization (Krueger et al, 2021) which encourages models to perform consistently across domains. In addition, self-supervised (Kim et al, 2021) and meta-learning-based techniques (Chen et al, 2022) have been explored to further improve generalization. However, without exposure to the target domain, generalization remains inherently limited.
2.2 Source-Free Domain Adaptation
Source-Free Domain Adaptation (SFDA) aims to adapt a pre-trained source model to an unlabeled target domain while ensuring that no source data is accessed during the adaptation process. By eliminating the dependence on source data, SFDA effectively safeguards source data privacy, making it particularly suitable for scenarios where data sharing is restricted. SFDA techniques can be broadly categorized into two main approaches: pseudo-label strategies and generative methods. The former leverages target pseudo-labels to facilitate self-training, thereby enabling implicit adaptation without requiring explicit supervision (Tanwisuth et al, 2021; Ahmed et al, 2021; Liang et al, 2021; Xin et al, 2023). The latter employs generative models to synthesize target-style training data, allowing the model to bridge the domain gap through data augmentation and distribution alignment (Qiu et al, 2021; Liu et al, 2021). Similar to SFDA, our proposed Few-Shot Test Time Adaptation (FS-TTA) also maintains the source-free property, ensuring that adaptation is performed without relying on source data while leveraging a small support set to enhance adaptation efficiency.
2.3 Few-Shot Transfer Learning
Test Time Adaptation (TTA) aims to adapt a pre-trained source model on-the-fly during inference to mitigate distribution shifts. Early TTA methods apply self-supervised learning objectives (Sun et al, 2020), but typically require access to training data or modification of the training process. TENT (Wang et al, 2021) addresses this by proposing fully test-time adaptation, relying solely on target samples and adapting batch normalization parameters via entropy minimization. Subsequent approaches such as (Schneider et al, 2020; Nado et al, 2020) update statistics on each incoming mini-batch, while methods like LAME (Boudiaf et al, 2022) and EATA (Niu et al, 2022) tackle catastrophic forgetting during continual adaptation. TSD (Wang et al, 2023) further integrates self-training to selectively update using confident predictions. As a result, adaptation often relies heavily on the quality of incoming test samples.
2.4 Test Time Adaptation
Test Time Adaptation (TTA) aims to adapt a pre-trained source model during inference to mitigate distribution shifts between training and test domains. Early TTA methods address this challenge through self-supervised auxiliary tasks (Sun et al, 2020), which, while effective, often require access to training data or modifications to the training procedure. To overcome this limitation, TENT (Wang et al, 2021) proposes fully test-time adaptation by leveraging only target data, updating batch normalization parameters via entropy minimization. Building on this, subsequent works (Schneider et al, 2020; Nado et al, 2020) estimate batch normalization statistics dynamically from incoming test batches. Other approaches, such as LAME (Boudiaf et al, 2022) and EATA (Niu et al, 2022), focus on preventing catastrophic forgetting during continuous model updates. More recently, TSD (Wang et al, 2023) incorporates self-training by selecting confident test samples to guide adaptation. Despite their progress, these methods heavily rely on the quality and stability of online target data.
2.5 Comparisons with Other Settings
@@
We compare Few-Shot Test Time Adaptation (FS-TTA) with similar problem settings (details are in the appendix), as illustrated in Table 1.
@@
  • Compared with Domain Generalization, FS-TTA eliminates the necessity of accessing source data, thereby ensuring the preservation of source data privacy. Moreover, it allows for adaptation to the downstream target domain by updating model parameters, making it more flexible and applicable in real-world settings.
  • Compared with Source-Free Domain Adaptation, FS-TTA removes the constraint of requiring all target domain data to be available at once. Instead, it facilitates dynamic and continuous online model updates, enabling adaptation based on incoming mini-batches of target data, which is particularly beneficial in streaming or real-time applications.
@@
  • Compared with Few-Shot Transfer Learning, FS-TTA not only makes use of a limited number of target domain samples for adaptation but also continuously refines the model during test time by incorporating online mini-batch target data. This ensures more efficient and progressive adaptation to changing target distributions.
@@
  • Compared with Test Time Adaptation, FS-TTA leverages a small auxiliary set of target samples, allowing the pre-trained source model to adapt more quickly and effectively to the target domain. Additionally, FS-TTA demonstrates superior performance in handling challenging scenarios where there are substantial domain shifts, making it a more robust and reliable solution.
3 Preliminary
3.1 Instance Normalization
Instance Normalization (Ulyanov et al, 2016) is a normalization technique widely used in deep neural network architectures, especially in the context of style transfer and generative models. Let us consider a batch images with size $N\times C\times H\times W$, where $N$ is the batch size, $C$ is the number of channels, and $H$ and $W$ are the height and width of the images. For each sample $i$ and channel $c$, we compute the mean and standard deviation as follows:
$\mu_{i,c}=\frac{1}{H\times W}\sum_{h=1}^{H}\sum_{w=1}^{W}x_{i,c,h,w},$ (1)
$\sigma_{i,c}=\sqrt{\frac{1}{H\times W}\sum_{h=1}^{H}\sum_{w=1}^{W}(x_{i,c,h,w}-\mu_{i,c})^{2}},$ (2)
where $x_{i,c,h,w}$ denotes the input feature of samples $i$, channel $c$, height $h$, and width $w$. After computing the mean and standard deviation, we can normalize the input features:
${\rm IN}(x_{i,c,h,w})=\gamma\frac{x_{i,c,h,w}-\mu_{i,c}}{\sigma_{i,c}}+\beta,$ (3)
where $\gamma,\beta\in R^{C}$ are learnable transformation parameters.
3.2 Class Prototype
The class prototype is a representative point in the feature space that summarizes the key characteristics of a class. For each class, it serves as a centroid or an anchor point around which samples of the class cluster. Let us denote $F=\{f_{1},f_{2},...,f_{n}\}$ as a set of $n$ sample features in class $c$, where each $f_{i}\in R^{d}$ represents a $d$-dimensional feature vector of a sample. The prototype $P$ of the class $c$ is calculated as the mean of all feature vectors, namely that:
$P_{c}=\frac{1}{n}\sum_{i=1}^{n}f_{i},$ (4)
where $P_{c}\in R^{d}$. The class prototype plays an important role in few-shot scenarios.
@@ 3.2 Class Prototype
4 Method
4.1 Problem Setting
Considering a typical scenario where a source model $f_{\theta_{s}}$ is equipped with parameters $\theta_{s}$ and trained on source datasets $\mathcal{D}{s_{1}},\mathcal{D}{s_{2}},…,\mathcal{D}{s_{n}}$, our objective is to adapt this pre-trained model to a target domain $D_{t}$ without accessing source data. A small, labeled support set $S={(s_{i},y_{i})}$ is provided from $D_{t}$, where $s_{i}$ denotes an image and $y_{i}$ its corresponding label. During test time, unlabeled target samples arrive sequentially in mini-batches. Few-Shot Test Time Adaptation (FS-TTA) aims to effectively adapt the source model $f_{\theta_{s}}$, by leveraging the support set $S$ in conjunction with the streaming unlabeled data to mitigate domain shift. Notably, the support set $S$ can be acquired offline prior to deployment, and in many real-world applications, collecting such limited supervision is both feasible and cost-efficient.
4.2 Stage I: Model Adaptation via Fine-Tuning
To significantly and swiftly enhance the initialization performance of the pre-trained source model in the target domain and minimize domain shifts, we design to fine-tune the pre-trained source model using the few-shot support set. Given the limited number of samples per class, there is a potential risk of overfitting during the fine-tuning process. To mitigate this, we introduce the Feature Diversity Augmentation (FDA) module, which generates new features by mixing statistics. Ultimately, we use a supervised classification loss to fine-tune the pre-trained source model. This entire procedure is illustrated in Stage I of Figure 3.
Feature Diversity Augmentation (FDA). Prior research (Zhou et al, 2021) has demonstrated a significant association between feature statistics and image style, which is intricately linked to data distribution within the field of computer vision. To increase style diversity while preserving semantic consistency, we introduce Feature Diversity Augmentation (FDA), a feature-level data augmentation technique that simulates various image styles without altering the original class labels. This approach effectively enriches the support set and helps reduce the risk of overfitting during fine-tuning.
FDA is incorporated between layers (blocks) in the pre-trained source backbone, as depicted in Figure 3. More specifically, FDA mixes the feature statistics of two random samples to generate new features. The computations within the FDA module can be summarized in three steps. Firstly, given two feature maps $f_{i}$ and $f_{j}$ from the support set, we compute their feature statistics ($\mu_{i},\sigma_{i}$) and ($\mu_{j},\sigma_{j}$). Secondly, FDA generates the mixtures of feature statistics:
$\bm{\gamma}_{\text{mix }}=\lambda\sigma_{i}+(1-\lambda)\sigma_{j},$ (5)
$\bm{\beta}_{\text{mix }}=\lambda\mu_{i}+(1-\lambda)\mu_{j}.$ (6)